The resistance of a bulb filament is 100 ? at a temperature of 100°C. If its temperature coefficient of resistance be 0.005 per °C, its resistance will become 200 ? at a temperature of ?

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
From the given data when applied to the formula R0=R(1+alpha∆t) we get
200 = 100[1 +(0.005 × ∆t)]
T− 100 = 200
T = 300° C
Thanks and Regards
Shaik Aasif
askIITians faculty
Ketan Prakash
5 years ago
When variation in temperature is more then, 
         R¹=R(1+alpha Δt) can not be used we have to use   R¹=R e ^(alphaΔt)
     Because we can not Neglect any effect on changing hind temperature.
           R¹/R= e^(alpha Δt)  
          ln(R¹/R)=alpha Δt 
    Hence the correct answer is 240 degree celsius.